OpenStudy (anonymous):

Find the value of Q in the following system so that the solution to the system is {(x,y) : x-3y-4} x-3y=4 2x-6y=Q

OpenStudy (anonymous):

so the solution is meant to be x-3y-4 and the two equations are x-3y=4 2x-6y=Q correct?

OpenStudy (anonymous):

sorry the first part should be {(x,y) : x-3y=4}

OpenStudy (anonymous):

The rest is right

OpenStudy (anonymous):

The equations are the same...the bottom one represents the top equation multiplied by 2

OpenStudy (anonymous):

If that's the case, what would Q be equal to?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Yes...that's what I 'd say

OpenStudy (anonymous):

Thank youu so much